OpenStudy (pd03922):

What is the value of the function y=3 - 1 when x= -1 A. -8 B. -4 C. 2 D. 6

OpenStudy (smartperson):

where is the x in the equation

OpenStudy (smartperson):

ohhhhhh i get it -1=x so the answer to y is x so it would be c

OpenStudy (pd03922):

Sorry it is supposed to look like this y= 3x - 1

OpenStudy (will.h):

Y = 3(-1) -1 Y = -3 -1 Y = -4

OpenStudy (pd03922):

Thanks!

OpenStudy (will.h):

No problem

OpenStudy (smartperson):

yeah he got it then